OpenStudy (anonymous):

Consider a circle with a radius of 5 centered at origin. Show that the tangent line at any point is perpendicular to the line segment joining the point with the center of the circle.

OpenStudy (phi):

what course is this?

OpenStudy (anonymous):

calc 1

OpenStudy (phi):

If you find the slope m of the tangent at point (x0,y0) on the circle, and the radius to point (x0,y0) has slope -1/m then you have proven it.

OpenStudy (anonymous):

how do we know it has that slope?

OpenStudy (phi):

use calculus? First step, the derivative of a curve gives the slope of a line tangent to the curve. find the derivative of x^2 + y^2 = 25 so you can find dy/dx

OpenStudy (anonymous):

2x+2y=0 ?

OpenStudy (phi):

treating both x and y as variables \[ \frac{d}{dx} x^2 = 2x \frac{dx}{dx}= 2x \] \[ \frac{d}{dx} y^2 = 2y \frac{dy}{dx} \]

OpenStudy (phi):

now solve for dy/dx (slope of the tangent)

OpenStudy (anonymous):

-x/y

OpenStudy (phi):

now if x0,y0 is a point on the circle with center 0,0 what is the slope of the line ?

OpenStudy (anonymous):

Would it be wrong to put in zero?

OpenStudy (phi):

I don't follow you.